OpenStudy (anonymous):

In certain experiments, the error made in determining the solubility of a substance is a random variable having the uniform density with a=-.025, beta = .025. What are the probabilities that such an error will be (a) between .010 and .015? (b)-.012 and .012?

OpenStudy (anonymous):

using the uniform distirubtion formula I get \[\frac{ 1 }{ \beta - \alpha }=\int\limits_{.010}^{.015}66.65x dx\] for the integral

OpenStudy (anonymous):

after plugging in the limits I get .3333 and 1.5996. Is this right?